I. You are just told by your doctor that you tested positive for a serious disease. The test has 99% accuracy, which means that the probability of testing positive given that you have the disease is 0.99, and also that the probability of testing negative given that you do not have the disease is 0.99. The good news is that this is a rare disease, striking only 1 in 10,000 people. a (5 points) Why is it good news that the disease is rare? b (10 points) What are the chances that you actually have the disease? Show your work.
Answer #1

a)as disease is very rare ; therefore conditional probability that the person have the disease given tested positive will be very less.Therefore this is a good news,

b)

P(tested positive)=P(have disease and tested positive)+P(not have disease and tested positive)

=(1/10000)*0.99+(1-1/10000)*(1-0.99)=0.010098

hence P(have the disease given tested positive)

=P(have disease and tested positive)/P(tested positive)=(1/10000)*0.99/0.010098=0.009804